In the Keynesian cross, assume that the consumption function is given by C=200+0.75(Y-T) Planned investment is 100; government purchases and taxes are both 100. a) Graph planned expenditure as a function of income. b) What is the equilibrium level of income? c) If government purchases increase to 125, what is the new equilibrium income? d) What level of government purchases is needed to achieve an income of 1,600?
Answer #1

Answer)

a) E=C+I+G=200+0.75(Y-100)+100+100

or
E=325+0.75Y

Slope= I Planned expend. 350 Slope .07.5 325 1,300 1,400

b) E=Y; Y=325+0.75Y

0.25Y = 325

Y=1,300

c)This effect can be solved using the multiplier.

ΔY = (1/1-0.75)ΔG

ΔY = Δ4*0.25 = 100

that is ,

Y' = Y + ΔY = 1300 + 100 = 1400

d) Y=200+0.75(Y-100)+100+G

1,600=200+0.75(1,600-100)+100+G

G=175